OpenStudy (anonymous):

solve by completing the square x^2 - 2x + 24

OpenStudy (anonymous):

x=1 plus or minus radical 5

OpenStudy (anonymous):

do you think you can explain to me how to do completing the square?

sam (.sam.):

(x-1)^2 +23

OpenStudy (anonymous):

^ wait nvm.. I messed up sorry.

sam (.sam.):

First you divide middle coefficient by 2, and then put (x-b)^2 +b^2 +C b is the middle number , then take b out and square it as shown. the C is the y-intercept or the one without any "x" multiplied by it

sam (.sam.):

o wait , in (x-b)^2, the b must be divided by 2

OpenStudy (anonymous):

x=1 plus or minus radical 23 Basically you move the 24 to the other side. So you have : x^2-2x=-24 now look at the 2x... all you have to do is divide the coefficient by two... so it'd be 2/2 which is one.. and then you square the 1... which is one... now that's that number you should add to both sides.. so it'd be: x^2-2x+1=-23 which can simplified into (x-1)^2=-23 next you get the squareroot of each side, and then isolate the x :)

OpenStudy (anonymous):

Oh okay i get that now :) thanks that cleared it up a lot better :)